### v3

# SynCo: Synthetic Hard Negatives in Contrastive Learning for Better Unsupervised Visual Representations

Abstract
Contrastive learning has become a dominant approach in self-supervised visual representation learning. Hard negatives - samples closely resembling the anchor - are key to enhancing learned representations’ discriminative power. However, efficiently leveraging hard negatives remains challenging. We introduce SynCo (Synthetic negatives in Contrastive learning), a novel approach that improves model performance by generating synthetic hard negatives on the representation space. Building on the MoCo framework, SynCo introduces six strategies for creating diverse synthetic hard negatives on-the-fly with minimal computational overhead. SynCo achieves faster training and better representation learning, reaching 67.9% top-1 accuracy on ImageNet ILSVRC-201 linear evaluation after 200 pretraining epochs, surpassing MoCo’s 67.5% using the same ResNet-50 encoder. It also transfers more effectively to detection tasks: on PASCAL VOC, it outperforms both the supervised baseline and MoCo with 82.6% AP; on COCO, it sets new benchmarks with 41.0% AP for bounding box detection and 35.7% AP for instance segmentation. Our synthetic hard negative generation approach significantly enhances visual representations learned through self-supervised contrastive learning. Code is available at https://github.com/giakoumoglou/synco.
1 Introduction
Figure 1: SynCo extends MoCo by introducing synthetic hard negatives generated on-the-fly from a memory queue. The process begins with two augmented views of an image, $\mathbf{x}_{q}$ and $\mathbf{x}_{k}$, processed by an encoder and a momentum encoder, respectively, producing feature vectors $\mathbf{q}$ and $\mathbf{k}$. The memory queue holds negative samples $\mathbf{n}_{1},\mathbf{n}_{2},\ldots$, which are concatenated with synthetic hard negatives $\mathbf{h}_{1},\mathbf{h}_{2},\ldots$ generated using the SynCo strategies. These combined negatives are used to compute the affinity matrix, which, together with the positive pair (query $\mathbf{q}$ and key $\mathbf{k}$), contributes to the InfoNCE loss calculation.
Contrastive learning has emerged as a prominent approach in self-supervised learning, significantly advancing representation learning from unlabeled data. This technique, which discriminates between similar and dissimilar data pairs, has shown premise in visual representation tasks. Seminal works such as SimCLR [6] and MoCo [20] established instance discrimination as a pretext task. These methods generate multiple views of the same data point through augmentation, training the model to minimize the distance between positive pairs (augmented views of the same instance) while maximizing it for negative pairs (views of different instances).
Despite its effectiveness, instance discrimination faces challenges. A key limitation is the need for numerous negative samples, often leading to increased computational costs. For example, SimCLR requires large batch sizes for sufficient negatives [6]. While approaches like MoCo address some issues through dynamic queues and momentum encoders [20, 8], they still face challenges in selecting and maintaining high-quality hard negatives.
Recent studies have highlighted the importance of carefully crafted data augmentations in learning robust representations [6, 11, 39, 45, 8, 23, 34, 3]. These transformations likely provide more diverse, challenging copies of images, increasing the difficulty of the self-supervised task. This self-supervised task is a pretext problem (e.g., predicting image rotations [13] or solving jigsaw puzzles [32]) designed to induce learning of generalizable features without explicit labels. Moreover, techniques that combine data at the pixel level [53, 50] or at the feature level [43] have proven effective in helping models learn more resilient features, leading to improvements in both fully supervised and semi-supervised tasks.
The concept of challenging negative samples has been explored as a way to enhance contrastive learning models. These samples, which lie close to the decision boundary, are crucial for refining the model’s discriminative abilities. Recent work like MoCHI [23] has shown improvements by incorporating harder negatives. Our work builds on this idea by proposing novel strategies for synthetic hard negative generation.
In this paper, we present SynCo (Synthetic negatives in Contrastive learning), a novel approach to contrastive learning that leverages synthetic hard negatives to enhance the learning process. Building on the foundations of MoCo, SynCo introduces six distinct strategies for generating synthetic hard negatives, each designed to provide diverse and challenging contrasts to the model. These strategies include: interpolated negatives; extrapolated negatives; mixup negatives; noise-injected negatives; perturbed negatives; and adversarial negatives. By incorporating these synthetic samples, SynCo aims to push the boundaries of contrastive learning, improving both the efficiency and effectiveness of the training process.
The main contributions of our work are as follows:
  1. 1.
    We introduce SynCo, a novel contrastive learning approach that improves representation learning by leveraging synthetic hard negatives. SynCo enhances the discriminative capabilities of models by generating challenging negatives on-the-fly from a memory queue, using six distinct strategies that target different aspects of the feature space to provide a comprehensive approach to hard negative generation. This process improves model performance without significant increases in computational demands, achieving faster training and better representation learning.
  2. 2.
    We empirically show improved downstream performance on ImageNet by incorporating synthetic hard negatives, demonstrating enhancements in both linear evaluation and semi-supervised learning tasks.
  3. 3.
    We show SynCo learns stronger representations by measuring their transfer learning capabilities COCO and PASCAL VOC detection, where it outperforms both the supervised baseline and MoCo.
The paper is structured as follows: Section 2 reviews related work; Section 3 explores hard negatives in contrastive learning; Section 4 introduces our synthetic hard negatives method; Section 5 presents experimental results; and Section 6 concludes the paper.
2 Related Work
2.1 Contrastive Learning
Recent contrastive learning methods focus on instance discrimination as a pretext task, treating each image as its own class. The core principle involves bringing an anchor and a ”positive” sample closer in the representation space while pushing the anchor away from ”negative” samples [6, 24]. Positive pairs are typically created through multiple views of each data point [39], using techniques such as color decomposition [38], random augmentation [6, 20], image patches [42], or student-teacher model representations [14, 5, 33]. The common training objective, based on InfoNCE [42] or its variants [6, 11, 41], aims to maximize mutual information [21, 2], necessitating numerous negative pairs. While some approaches use large batch sizes [6] to address this, others like MoCo [20, 8], PIRL [30], and InstDis [47] employ memory structures. Recent advancements explore strategies such as invariance regularizers [31], dataset-derived positives [11], and unified contrastive formulas [37]. Some methods eliminate negative samples through asymmetric Siamese structures or normalization [14, 7, 5], while others prevent model collapse via redundancy reduction [51] or regularization [4]. Despite these innovations, the fine-grained nature of instance discrimination can lead to false-negative pairs [55].
2.2 Hard Negatives
Hard negatives are critical in contrastive learning as they improve the quality of visual representations by helping to define the representation space more effectively. These challenging yet relevant samples are harder to distinguish from the anchor point, enabling the model to better differentiate between similar features. The use of hard negatives involves selecting samples that are similar to positive samples but different enough to aid in learning distinctive features. Dynamic sampling of hard negatives during training prevents the model from easily minimizing the loss, enhancing its learning capabilities [20, 6]. Various approaches have been proposed to leverage hard negatives effectively. For instance, MoCo [20] utilizes a dynamic queue and momentum-based encoder updates to maintain fresh and challenging negatives throughout training. Other methods, such as SimCLR [6] and InfoMin [39], suggest adjusting the difficulty of negative samples by varying data augmentation techniques. This progressive increase in task difficulty benefits the training process. Building on these ideas, MoCHI [23] has explored integrating hard negative mixing into existing frameworks to further improve performance. When incorporating these approaches into representation space generation, it’s essential to consider how negative samples impact the model’s learning space. The goal is to implement a strategy that not only challenges the model but also enhances its ability to extract and differentiate features effectively. By employing these methods, models become more adept at handling detailed and complex tasks, ensuring each negative sample significantly contributes to optimizing learning outcomes and boosting overall model effectiveness.
2.3 Synthetic Features
Synthetic feature generation is a widely used method to enhance deep learning models, especially in cases with limited labeled data. By adding synthetic features to the representation space, models improve in generalization and performance. Methods have generated features for unseen classes using generative models [17, 48, 36], while others have integrated these into self-supervised and contrastive learning frameworks [27, 54]. Models that combine generation with representation spaces have also shown success in zero-shot learning [16]. In contrast, our approach directly generates synthetic hard negatives in contrastive learning, improving representation without additional generative models.
3 Preliminaries
3.1 Contrastive Learning
Contrastive learning seeks to differentiate between similar and dissimilar data pairs, often treated as a dictionary look-up where representations are optimized to align positively paired data through contrastive loss in the representation space [20]. Given an image $x$, and a distribution of image augmentation $\mathcal{T}$, we create two augmented views of the same image using the transformation $t_{q},t_{k}\sim\mathcal{T}$, i.e., $x_{q}=t_{q}(x)$ and $x_{k}=t_{k}(x)$. Two encoders, $f_{q}$ and $f_{k}$, namely the query and key encoders, generate the vectors $\mathbf{q}=f_{q}(x_{q})$ and $\mathbf{k}=f_{k}(x_{k})$ respectively. The learning objective minimizes a contrastive loss using the InfoNCE criterion [42]:
$\mathcal{L}(\mathbf{q},\mathbf{k},\mathcal{Q})=-\log\frac{\exp(\mathbf{q}^{T}\cdot\mathbf{k}/\tau)}{\exp(\mathbf{q}^{T}\cdot\mathbf{k}/\tau)+\sum\limits_{\mathbf{n}\in\mathcal{Q}}\exp(\mathbf{q}^{T}\cdot\mathbf{n}/\tau)}$ (1)
Here, $\mathbf{k}$ is $f_{k}$’s output from the same augmented image as $\mathbf{q}$, and $\mathcal{Q}=\{\mathbf{n}_{1},\mathbf{n}_{2},\ldots,\mathbf{n}_{K}\}$ includes outputs from different images, representing negative samples of size $K$. The temperature parameter $\tau$ adjusts scaling for the $\ell_{2}$-normalized vectors $\mathbf{q}$ and $\mathbf{k}$. The key encoder $f_{k}$ can be updated in two ways. In the synchronized update approach, $f_{k}$ is updated synchronously with $f_{q}$, maintaining identical weights throughout training [6]. Alternatively, a momentum update scheme can be employed, where $f_{k}$ is updated using the equation: $\theta_{k}\leftarrow m\cdot\theta_{k}+(1-m)\cdot\theta_{q}$ [20]. Here, $\theta_{k}$ and $\theta_{q}$ are the parameters of $f_{k}$ and $f_{q}$ respectively, and $m\in[0,1]$ is the momentum coefficient. This momentum approach allows $f_{k}$ to evolve more slowly, providing more consistent negative samples over time and potentially stabilizing the learning process. The memory bank $\mathcal{Q}$ can be defined in various ways, such as an external memory of all dataset images [30, 38, 47], a queue of recent batches [20], or the current minibatch [6].
The gradient of the contrastive loss in Equation 1 with respect to the query $\mathbf{q}$ is given by:
$\frac{\partial\mathcal{L}(\mathbf{q},\mathbf{k},\mathcal{Q})}{\partial\mathbf{q}}=-\frac{1}{\tau}\left((1-p_{k})\cdot\mathbf{k}-\sum_{\mathbf{n}\in\mathcal{Q}}p_{n}\cdot\mathbf{n}\right)$ (2)
where
$p_{z_{i}}=\frac{\exp(\mathbf{q}^{T}\cdot\mathbf{z_{i}}/\tau)}{\sum_{j\in Z}\exp(\mathbf{q}^{T}\cdot\mathbf{z_{j}}/\tau)}$ (3)
with $\mathbf{z_{i}}$ being a member of the set $\mathcal{Q}\cup\{\mathbf{k}\}$. The positive and negative logits contribute to the loss similarly to a $(K+1)$-way cross-entropy classification, with the key logit representing the query’s latent class [1].
3.2 Understanding Hard Negatives
The effectiveness of contrastive learning approaches hinges critically on the utilization of hard negatives [1, 15, 22, 29, 47, 23]. Current approaches face significant challenges in efficiently leveraging these hard negatives. Sampling from within the same batch necessitates larger batch sizes [6, 9], potentially straining computational resources. Conversely, maintaining a memory bank containing representations of the entire dataset incurs substantial computational overhead in keeping the memory up-to-date [30, 47, 20, 8]. These limitations underscore the need for more efficient strategies to generate and utilize hard negatives in contrastive learning frameworks.
Hardness of negatives. The ”hardness” of negative samples, defined by their similarity to positive samples in the representation space, determines how challenging they are for the model to differentiate, directly impacting the effectiveness of the contrastive learning process. Figure 2 illustrates the evolution of negative sample hardness during MoCo-v2 training on ImageNet-100. The plot depicts the top 1024 matching probabilities $p_{z_{i}}$ across different training epochs. Initially, the distribution of these probabilities is relatively uniform. However, as training progresses, a clear trend emerges: fewer negatives contribute significantly to the loss function. This observation suggests that the model rapidly learns to distinguish most negatives, leaving only a small subset that remains challenging. Such a phenomenon underscores the importance of maintaining a diverse pool of hard negatives throughout the training process to sustain effective learning [23].
Figure 2: Histogram of the top 1024 matching probabilities $p_{z_{i}}$, $z_{i}\in\mathcal{Q}$ for MoCo-v2, over various epochs. Logits are organized in descending order, and each line indicates the mean matching probability across all queries [23].
Difficulty of the proxy task. The difficulty of the proxy task in contrastive learning, typically defined by the self-supervised objective, significantly influences the quality of learned representations. Figure 3 compares the proxy task performance of MoCo and MoCo-v2 on ImageNet-100, measured by the percentage of queries where the key ranks above all negatives. Notably, MoCo-v2, which employs more aggressive augmentations, exhibits lower proxy task performance compared to MoCo, indicating a more challenging learning objective. Paradoxically, this increased difficulty correlates with improved performance on downstream tasks such as linear classification [23]. This counterintuitive relationship between proxy task difficulty and downstream performance suggests that more challenging self-supervised objectives can lead to the learning of more robust and transferable representations, motivating the development of strategies to dynamically modulate task difficulty during training.
Figure 3: Performance comparison of MoCo, MoCo-v2, MoCHI, and SynCo (under various configurations) on ImageNet-100 in terms of accuracy on the proxy task (percentage of queries where the key is ranked higher than all negatives).
4 Synthetic Hard Negatives in Contrastive Learning
Based on these observations, we present an approach for generating synthetic hard negatives in the representation space using six distinct strategies. We refer to the proposed approach as SynCo (”Synthetic negatives in Contrastive learning”).
4.1 Hard Negatives Generation
Let $\mathbf{q}$ represent the query image, $\mathbf{k}$ its corresponding key, and $\mathbf{n}\in\mathcal{Q}$ denote the negative features from a memory structure of size $K$. The loss associated with the query is computed using the logits $\ell(\mathbf{z_{i}})=\mathbf{q}^{T}\cdot\mathbf{z_{i}}/\tau$, which are processed through a softmax function. We define $\mathcal{\hat{Q}}=\{\mathbf{n}_{1},\mathbf{n}_{2},\ldots,\mathbf{n}_{K}\}$ as the ordered set of all negative features, where $\ell(\mathbf{n}_{i})>\ell(\mathbf{n}_{j})$ for all $i<j$, implying that the negative features are sorted based on decreasing similarity to the query. The most challenging negatives are selected by truncating the ordered set $\mathcal{\hat{Q}}$, retaining only the first $N<K$ elements, denoted as $\mathcal{\hat{Q}}^{N}$.
Interpolated hard negatives (type 1). For each query $\mathbf{q}$, we propose to generate $N_{1}$ synthetic hard negative features by mixing the query $\mathbf{q}$ with a randomly chosen feature from the $N$ hardest negatives in $\mathcal{\hat{Q}}^{N}$. Let $H^{1}=\{\mathbf{h}_{1}^{1},\mathbf{h}_{2}^{1},\ldots,\mathbf{h}_{N_{1}}^{1}\}$ be the set of synthetic negatives to be generated. Then a synthetic negative feature $\mathbf{h}_{k}^{1}\in H^{1}$ would be given by:
$\mathbf{h}_{k}^{1}=\alpha_{k}\cdot\mathbf{q}+(1-\alpha_{k})\cdot\mathbf{n}_{i},\quad\alpha_{k}\in(0,\alpha_{\text{max}})$ (4)
where $\mathbf{n}_{i}\in\mathcal{\hat{Q}}^{N}$ and $\alpha_{k}$ is randomly sampled from a uniform distribution in the range $(0,\alpha_{\text{max}})$. The resulting synthetic hard negatives are then normalized and added to the set of negative logits for the query. Interpolation creates a synthetic embedding that lies between the query and the negative in the representation space. We set $\alpha_{\text{max}}=0.5$ to guarantee that the contribution of the query is always less than that of the negative. This is similar to the hardest negatives (type 2) of MoCHI [23].
Extrapolated hard negatives (type 2). For each query $\mathbf{q}$, we propose to generate $N_{2}$ hard negative features by extrapolating beyond the query embedding in the direction of the hardest negative features. Similar to the interpolated method, we use a randomly chosen feature from the $N$ hardest negatives in $\mathcal{\hat{Q}}^{N}$. Let $H^{2}=\{\mathbf{h}_{1}^{2},\mathbf{h}_{2}^{2},\ldots,\mathbf{h}_{N_{2}}^{2}\}$ be the set of synthetic negatives to be generated. Then a synthetic negative feature $\mathbf{h}_{k}^{2}\in H^{2}$ would be given by:
$\mathbf{h}_{k}^{2}=\mathbf{n}_{i}+\beta_{k}\cdot(\mathbf{n}_{i}-\mathbf{q}),\quad\beta_{k}\in(1,\beta_{\text{max}})$ (5)
where $\mathbf{n}_{i}\in\mathcal{\hat{Q}}^{N}$ and $\beta_{k}$ is randomly sampled from a uniform distribution in the range $(1,\beta_{\text{max}})$. These synthetic features are also normalized and used to enhance the negative logits. Extrapolation generates a synthetic embedding that lies beyond the query embedding in the direction of the hardest negative. We choose $\beta_{\text{max}}=1.5$.
Mixup hard negatives (type 3). For each query $\mathbf{q}$, we propose to generate $N_{3}$ hard negative features by combining pairs of the $N$ hardest existing negative features in $\mathcal{\hat{Q}}^{N}$. Let $H^{3}=\{\mathbf{h}_{1}^{3},\mathbf{h}_{2}^{3},\ldots,\mathbf{h}_{N_{3}}^{3}\}$ be the set of synthetic negatives to be generated. Then a synthetic negative feature $\mathbf{h}_{k}^{3}\in H^{3}$ would be given by:
$\mathbf{h}_{k}^{3}=\gamma_{k}\cdot\mathbf{n}_{i}+(1-\gamma_{k})\cdot\mathbf{n}_{j},\quad\gamma_{k}\in(0,1)$ (6)
where $\mathbf{n}_{i},\mathbf{n}_{j}\in\mathcal{\hat{Q}}^{N}$ and $\gamma_{k}$ is randomly sampled from a uniform distribution in the range $(0,1)$. The resulting synthetic hard negatives are then normalized and added to the set of negative logits for the query. Mixup combines pairs of the hardest existing negative features to create a synthetic embedding that represents a blend of challenging cases. This is similar to the hard negatives (type 1) of MoCHI [23].
Noise-injected hard negatives (type 4). For each query $\mathbf{q}$, we propose to generate $N_{4}$ hard negative features by adding Gaussian noise to the hardest negative features. Using the top $N$ hardest negatives $\mathcal{\hat{Q}}^{N}$, let $H^{4}=\{\mathbf{h}_{1}^{4},\mathbf{h}_{2}^{4},\ldots,\mathbf{h}_{N_{4}}^{4}\}$ be the set of synthetic negatives to be generated. Then a synthetic negative feature $\mathbf{h}_{k}^{4}\in H^{4}$ would be given by:
$\mathbf{h}_{k}^{4}=\mathbf{n}_{i}+\mathcal{N}(\mathbf{0},\sigma^{2}\cdot\mathbf{I})$ (7)
where $\mathbf{n}_{i}\in\mathcal{\hat{Q}}^{N}$ and $\mathcal{N}(\mathbf{0},\sigma^{2}\cdot\mathbf{I})$ represents Gaussian noise with standard deviation $\sigma$ (where $\mathbf{I}$ is the identity matrix). The noisy negatives are normalized before being used in the loss calculation. Noise injection adds Gaussian noise to the hardest negative features, resulting in a synthetic embedding with added randomness.
Perturbed hard negatives (type 5). For each query $\mathbf{q}$, we propose to generate $N_{5}$ hard negative features by perturbing the embeddings of the hardest negative features. Given the top $N$ hardest negatives $\mathcal{\hat{Q}}^{N}$, let $H^{5}=\{\mathbf{h}_{1}^{5},\mathbf{h}_{2}^{5},\ldots,\mathbf{h}_{N_{5}}^{5}\}$ be the set of synthetic negatives to be generated. Then a synthetic negative feature $\mathbf{h}_{k}^{5}\in H^{5}$ would be given by:
$\mathbf{h}_{k}^{5}=\mathbf{n}_{i}+\delta\cdot\nabla_{\mathbf{n}_{i}}\text{sim}(\mathbf{q},\mathbf{n}_{i})$ (8)
where $\mathbf{n}_{i}\in\mathcal{\hat{Q}}^{N}$ and $\text{sim}(\cdot,\cdot)$ is the similarity function and $\delta$ controls the perturbation magnitude. The perturbed embeddings are then normalized and added to the negative logits. Perturbation modifies the embeddings of the hardest negative features based on the gradient of the similarity function, creating synthetic negatives that are slightly adjusted to be more challenging for the model.
Adversarial hard negatives (type 6). For each query $\mathbf{q}$, we propose to generate $N_{6}$ hard negative features by applying adversarial perturbations to the hardest negative features to maximize their similarity to the query embeddings. Using the top $N$ hardest negatives $\mathcal{\hat{Q}}^{N}$, let $H^{6}=\{\mathbf{h}_{1}^{6},\mathbf{h}_{2}^{6},\ldots,\mathbf{h}_{N_{6}}^{6}\}$ be the set of synthetic negatives to be generated. Then a synthetic negative feature $\mathbf{h}_{k}^{6}\in H^{6}$ would be given by:
$\mathbf{h}_{k}^{6}=\mathbf{n}_{i}+\eta\cdot\text{sign}(\nabla_{\mathbf{n}_{i}}\text{sim}(\mathbf{q},\mathbf{n}_{i}))$ (9)
where $\mathbf{n}_{i}\in\mathcal{\hat{Q}}^{N}$ and $\eta$ controls the perturbation magnitude. The perturbed embeddings are normalized and added to the negative logits. Adversarial hard negatives apply adversarial perturbations to the hardest negative features, specifically altering them to maximize their similarity to the query embeddings, thereby producing the most challenging contrasts.
4.2 Discussion on Synthetic Hard Negatives
Intuition of synthetic hard negatives. Each SynCo strategy imrpoves model generalization through challenging contrasts. Type 1 interpolates between query and hard negatives, increasing sample diversity throughout training. Type 2 extrapolates beyond the query, pushing representation space boundaries and improving robustness to difficult contrasts. Type 3 combines pairs of hard negatives, encouraging more generalized and robust feature learning. Type 4 injects Gaussian noise, promoting invariance to minor feature fluctuations and enhancing generalization. Type 5 modifies embeddings based on similarity gradients, refining discriminatory power by directing the model towards harder negatives. Type 6 applies adversarial perturbations, creating the most challenging contrasts to distinguish deceptively similar samples. A simplified example illustrating the synthetic hard negative generation is provided in Figure 4. The latter figure visualizes the different types of synthetic hard negatives generated from a query and three original negative features in a 2D representation space. Dotted lines connect the query to extrapolated negatives and between pairs of mixup negatives, illustrating the relationships used to generate these synthetic negative features.
Figure 4: A toy example illustrating the six types of synthetic hard negatives generated by SynCo from a query and three original negative features in a 2D representation space.
Is the proxy task more difficult? Figure 3 depicts the proxy task performance for different configurations of SynCo. We observe that incorporating synthetic negatives leads to faster learning and improved performance. Each type of synthetic negative accelerates learning compared to the MoCo-v2 baseline, with the full SynCo configuration showing the most significant improvement (see Figure 6(a)) and the lowest final proxy task performance. This indicates that SynCo presents the most challenging proxy task. This is evidenced by $\max\ell(\mathbf{h}_{k}^{i})>\max\ell(\mathbf{n}_{j})$, where $\mathbf{h}_{k}^{i}\in H^{i}$ are synthetic negatives and $\mathbf{n}_{j}\in\tilde{\mathcal{Q}}_{N}$ are original negatives. Through SynCo, we modulate proxy task difficulty via synthetic negatives, pushing the model to learn more robust features.
Evaluating the usage of the representation space. To assess learned representations, we employ alignment and uniformity metrics proposed by [44]. These metrics provide insights into representation space utilization, with alignment quantifying the grouping of similar samples and uniformity measuring representation spread across the hypersphere. Figure 5 presents results for various models using features from the ImageNet-100 validation set. Our findings demonstrate that SynCo significantly enhances the uniformity of representations compared to MoCo-v2 and MoCHI, indicating its superior ability to utilize the representation space in the proxy task. Furthermore, the incorporation of synthetic negatives (types 1 to 6) leads to improved alignment. These results suggest that SynCo’s approach to synthetic negative generation and contrastive learning yields more effective and well-distributed feature representations.
Figure 5: Performance comparison of MoCo-v2, MoCHI, and SynCo (under various configurations) on ImageNet-100 in terms of alignment and uniformity metrics. The x-axis and y-axis represent $-\mathcal{L}_{\text{uniform}}$ and $-\mathcal{L}_{\text{align}}$, respectively. The model with the highest performance is located in the upper-right corner of the chart. We use $K=65$k.
5 Experiments
5.1 Implementation Details
We pretrain SynCo on ImageNet ILSVRC-2012 [10] and its smaller ImageNet-100 subset [38] for ablation studies using a ResNet-50 [18] encoder. Our implementation builds upon MoCo-v2 [8]. For training, unless stated otherwise, we use $K=65$k ($K=16$k for ImageNet-100). For SynCo, we also have a warm-up of 10 epochs, i.e. for the first epochs we do not synthesize hard negatives. We set SynCo’s hyperparameters $\sigma$, $\delta$, and $\eta$ to 0.01. For hard negative generation, we use the top $N=1024$ hardest negatives, with $N_{1}=N_{2}=N_{3}=256$ and $N_{4}=N_{5}=N_{6}=64$. For ImageNet linear evaluation, we train a linear classifier on frozen features for 100 epochs, using a batch size of 256 and a cosine learning rate schedule. Initial learning rates are set to 30.0 for ImageNet and 10.0 for ImageNet-100. To evaluate transfer learning, we apply SynCo to object detection tasks. For PASCAL VOC [12], we fine-tune a Faster R-CNN [35] (R50-C4) on trainval07+12 and test on test2007. For COCO [28], we use a Mask R-CNN [19] (R50-C4), fine-tuning on train2017 and evaluating on val2017. We employ Detectron2 [46] and report standard AP metrics, following [20] without additional hyperparameter tuning. Detailed implementation details are provided in the supplementary material.
5.2 Linear Evaluation on ImageNet
We evaluate the SynCo representation by training a linear classifier on top of the frozen features that were pretrained on ImageNet, following the procedure described in [26, 25, 6, 14, 42] (details in supplementary material). SynCo obtains $67.9\%\pm 0.16\%$ top-1 accuracy and $88.0\%\pm 0.05\%$ top-5 accuracy after 200 epoch pretraining (Table 1). Specifically, SynCo achieves a +0.4% top-1 accuracy improvement over MoCo-v2 and a +1.0% improvement over MoCHI. While MoCHI, which employs hard negatives, achieves lower performance than MoCo-v2, our method, which generates synthetic hard negatives, not only avoids this drop in performance but actually improves it. SynCo also surpasses the state-of-the-art methods SimCLR and SimSiam. When training for 800 epochs, SynCo obtains 70.6% top-1 accuracy (89.8% top-5 accuracy) (Table 2). This represents a +1.9% improvement in top-1 accuracy over MoCHI. These results show that SynCo achieves higher accuracy than its competitors with only a minor computational overhead for generating synthetic hard negatives, while also enabling faster training convergence.
Table 1: Top-1 and top-5 accuracies (in %) under linear evaluation on ImageNet ILSVRC-2012 with 200 epochs of pretraining using ResNet-50. Results for SynCo are averaged over 3 runs.
Method Top-1 Top-5
Supervised 76.5 -
PIRL 63.6 -
LA 60.2 -
InfoMin 70.1 89.4
SimSiam 68.1 -
SimCLR 69.3 -
MoCo 60.7 -
MoCo-v2 67.5 90.1
MoCHI 66.9 -
SynCo (ours) 67.9 88.0
Table 2: Top-1 and top-5 accuracies (in %) under linear evaluation on ImageNet ILSVRC-2012 for models trained with extended epochs using ResNet-50. Results for SynCo are based on 1 run.
Method Epochs Top-1 Top-5
PIRL 800 63.6 -
InfoMin 800 73.0 91.1
BYOL 1000 74.3 91.6
SimSiam 800 68.1 -
SimCLR 1000 69.3 -
Barlow Twins 1000 73.2 91.0
MoCo-v2 800 71.1 90.1
MoCHI 800 68.7 -
SynCo (ours) 800 70.6 89.8
5.3 Semi-supervised Training on ImageNet
We evaluate SynCo in a semi-supervised setting using 1% and 10% of labeled ImageNet data, following protocols from [26, 52, 6, 14, 42] (details in supplementary material). We use the same data splits as [6]. Results in Table 3 show SynCo’s competitive performance. With 1% labeled data, SynCo achieves $50.8\%\pm 0.21\%$ top-1 accuracy and $77.5\%\pm 0.12\%$ top-5 accuracy. Using 10% labeled data, it reaches $66.6\%\pm 0.19\%$ top-1 accuracy and $88.0\%\pm 0.10\%$ top-5 accuracy. We observe that despite a lower volume of labeled data, SynCo competes well with established semi-supervised methods like SimCLR, BYOL, and Barlow Twins.
Table 3: Semi-supervised learning on ImageNet ILSVRC-2012 with 1% and 10% training examples using ResNet-50. Results for SynCo are averaged over 3 runs.
Method Epochs Top-1 Top-5
1% 10% 1% 10%
Supervised 25.4 56.4 48.4 80.4
InstDis 200 - - 39.2 77.4
PIRL 800 30.7 60.4 57.2 83.8
SimCLR 1000 48.3 65.6 75.5 87.8
Barlow Twins 1000 55.0 69.7 79.2 89.3
BYOL 1000 53.2 68.8 78.4 89.0
SwAV 800 53.9 70.2 78.5 89.9
SynCo (ours) 800 50.8 66.6 77.5 88.0
5.4 Transferring to Detection
We evaluate the SynCo representation, pretrained for 200 epochs, by applying it to the detection task. We follow the protocol of [20]. Detailed configurations for object detection experiments are provided in in the supplementary material. Table 4 presents the results of object detection on PASCAL VOC and COCO datasets. SynCo displays consistent gains over both the supervised baseline and MoCo-v2. Specifically, SynCo, surpasses MoCo-v2 and performs on par with MoCHI in the PASCAL VOC detection task. On the more challenging COCO dataset, SynCo achieves new state-of-the-art results in both bounding box detection and instance segmentation. Additional results comparing SynCo and MoCo-v2 at different training epochs are presented in the supplementary material.
Table 4: Results for object detection on PASCAL VOC and COCO. All models use a ResNet-50 backbone. We report AP, AP50, and AP75, which are standard COCO metrics. $bb$ denotes bounding box detection, and $msk$ denotes instance segmentation. The values in bold indicate the maximum of each column. Results for SynCo are averaged over 3 runs.
Method VOC07+12 detection COCO detection COCO segmentation
$AP$ $AP_{50}$ $AP_{75}$ $AP^{bb}$ $AP^{bb}_{50}$ $AP^{bb}_{75}$ $AP^{msk}$ $AP^{msk}_{50}$ $AP^{msk}_{75}$
Random init 16.8 35.9 13.0 31.0 49.5 33.2 28.5 46.8 30.4
Supervised 53.5 81.3 58.8 38.2 58.2 41.6 33.3 54.7 35.2
InstDis - - - 37.4 57.6 40.6 34.1 54.6 36.4
PIRL 55.5 81.0 61.3 38.5 57.6 41.2 34.0 54.6 36.2
InfoMin 57.6 82.7 64.6 39.0 58.5 42.0 34.1 55.2 36.3
SwAV 56.1 82.6 62.7 38.2 58.2 41.3 33.8 55.2 35.9
SimSiam 57.0 82.4 63.7 39.2 59.3 42.1 34.4 56.0 36.7
Barlow Twins 56.8 82.6 63.4 39.2 59.0 42.5 34.3 56.0 36.5
MoCo 55.5 81.5 61.3 38.5 58.3 41.6 33.6 54.8 35.6
MoCo-v2 57.0 82.4 63.6 39.0 58.6 41.9 34.2 55.4 36.2
MoCHI 57.5 82.7 64.4 39.2 58.9 42.4 34.4 55.5 36.6
SynCo (ours) 57.2 82.6 63.9 41.0 60.6 44.8 35.7 57.4 38.1
5.5 Ablation Study
Figure 6: Ablation studies on ImageNet-100 in terms of Top-1 and Top-5 accuracies (in %). (a) Performance comparison of different types of hard negatives. (b) Performance comparison of different values for $\sigma$, $\delta$, and $\eta$ on SynCo’s type 4, type 5, and type 6 hard negatives, respectively. (c) Comparison of SynCo and MoCo-v2 across different queue sizes. For all ablations we use $K=65$k.
We perform ablations studies on ImageNet-100. The results of our ablations are presented in Figure 6. Our findings consistently demonstrate that various SynCo configurations outperform the MoCo-v2 baseline. Additional ablation studies and analyses are presented in the supplementary material.
Type of hard negative. We evaluate the impact of each synthetic hard negative type on pretraining. For this, we select the top $N=1024$ hardest negatives and generate $N_{i}=256$, $i=1,2,\ldots,6$ negatives. We train SynCo without hard negatives (equivalent to MoCo-v2) for 100 epochs and measure top-1 and top-5 accuracy. Subsequently, we train SynCo using each type of hard negative individually, and then using all six types in combination. The results of these ablations are presented in Figure 6(a). We see that every SynCo configuration outperform the MoCo-v2 baseline.
Parameter sensitivity analysis. We conducted ablations on the parameters $\sigma$, $\delta$, and $\eta$ of SynCo’s type 4, type 5, and type 6 negatives, respectively. The results, presented in Figure 6(b), show that varying these parameters does not lead to significant differences in performance. This suggests that SynCo is robust across a wide range of values for $\sigma$, $\delta$, $\eta$.
Queue size. We investigate the effect of queue size $\mathcal{Q}$ on performance. We train SynCo and MoCo-v2 with reduced queue sizes. Our results, presented in Figure 6(c), reveal that SynCo performs comparably to MoCo-v2 across various queue sizes. With smaller queues, SynCo underperforms compared to MoCo-v2. This can be attributed to the fact that the total generated synthetic negatives are too hard for the task and harm performance, a finding that is also observed in [23]. However, as the queue increases, SynCo performs on par with MoCo-v2. At the largest queue size tested, SynCo outperforms MoCo-v2.
6 Conclusion
This paper introduces SynCo, a novel approach to contrastive learning that leverages synthetic hard negatives to enhance visual representation learning. By generating diverse and challenging negative samples on-the-fly, SynCo addresses the limitations of existing methods in maintaining a pool of effective hard negatives throughout training. Our comprehensive experiments demonstrate that SynCo not only accelerates the learning process but also leads to more robust and transferable representations. The effectiveness of SynCo is evidenced by its superior performance across various benchmarks, including linear evaluation on ImageNet, semi-supervised learning tasks, and transfer learning to object detection on PASCAL VOC and COCO datasets.
While our experiments primarily utilized the MoCo framework, the proposed hard negative generation strategies are general and can be applied to any contrastive learning method that benefits from hard negatives, including SimCLR[6], CPC [42], PIRL [30], and others [49, 45, 24, 11, 40]. These methods, which employ the InfoNCE loss function (or its variants [6, 11]) and instance discrimination as the pretext task, can benefit from the enhanced hard negative generation strategies proposed by SynCo. The introduction of synthetic hard negatives can provide these methods with more challenging, informative contrasts, potentially leading to better feature representations. Also, SynCo’s applicability extends beyond visual representation learning. The concept of hard negatives can be applied to various domains such as natural language processing, audio processing, and other areas where contrastive learning is relevant.
Broader Impact
The presented research should be categorized as research in the field of unsupervised learning. This work may inspire new algorithms, theoretical, and experimental investigation. The algorithm presented here can be used for many different vision applications and a particular use may have both positive or negative impacts, which is known as the dual use problem. Besides, as vision datasets could be biased, the representation learned by SynCo could be susceptible to replicate these biases.